) e we fail reject he null hypothesis at 5% significance. There is extrem he mean list price of four bedroom homes. hy strong evidence of a significance difference between the mean st price of three e t a homs nd 0 None of the above Question 19 A national computer retailer believes that the average sales are greater for salespersons with a college degree. A random sample of 3I salespersons with a degre had an avernge weekly sale of $3422 last year, while 35 salespersons without a college degree averaged $3255 in weekly sales. The standard devistions were $468 and $642 respectively. Is there evidence at the 5% level to support the retailers belief? Select the (p-value, Decision to Reject (RHo) or Failure to Reject (FRHO)1 a) Ip-value-0.015, RHol b) [p-value-0.117, FRHol c) O1p-value= 0.015, FRHo] d) o lp-value 0.117, RHo e) Ip-value-0.038, RHol Question 20 The community hospital is studying its distribution of patiensts. A random sample of 329 patbents presently in the hospital gave the following mformation Type of Patient old Rate of Occutrences Present Namber of Occurrences Maternity Ward 20% 79 95

please provide detailed explanation. I have a good grasp of the concept but I'm just trying to see where I went wrong on this problem.

I was able to plug in everything but my pvalue is incorrect.

If possible can someone include the rstudio code to do this problem.
Answer #1

The test statistic t = (\bar x1 - \bar x2)/sqrt(s1^2/n1 + s2^2/n2)

= (3422 - 3255)/sqrt((468)^2/31 + (642)^2/35) = 1.217

DF = ((468)^2/31 + (642)^2/35)^2/(((468)^2/31)^2/30 + ((642)^2/35)^2/34) = 62

P-VALUE = P(T > 1.217)

= 1 - P(T < 1.217)

= 1 - 0.8859 = 0.1141

As the p-value is greater than the significance level (0.1135 > 0.05), so the null hypothesis is not rejected.

Option-B is correct.